OpenStudy (anonymous):

Which equation is in point slope form? Point (7,4) Slope 6 a. y-7=6(x+4) b.y-4=6(x-7) c. y-4=6(x+7) d.y+4=6(x-7) Please help!!!!

jigglypuff314 (jigglypuff314):

Hello @buggie2323 and Welcome to OpenStudy! :) point slope form is: \(y - k = m(x-h)\) where we are given a point \((h,~k)\) and the slope \(m\)

jigglypuff314 (jigglypuff314):

In this case, the point = \((h,~k) = (7,~4)\) so we can plug in \(7\) for \(h\) and \(4\) for \(k\) also, we were given slope = \(6\) = \(m\) so we can plug in \(6\) for \(m\) all into the equation \(y - k = m(x - h)\) ^4 ^6 ^7 and just remember to keep the minuses I gave (don't change them) That should get you to your answer :) If you still need help with this question when you come back, feel free to tag me with @jigglypuff314 ^_^